OpenStudy (anonymous):

Does the table represent a linear or an exponential function? Explain. x 1 2 3 4 y 16 64 256 1024

OpenStudy (amistre64):

linear would indicate that there is a common value that can be added to each term to get the next

OpenStudy (amistre64):

16+16 = 64 64+16 not= 256 ; so it aint linear

OpenStudy (amistre64):

ack, 16+16 = 64 aint right, but it still aint linear :)